# Kernel Search approach to solve the Minimum Spanning Tree Problem with conflicting edge pairs00footnotetext: Email addresses:

### v1

@@ 4 Computational Tests
In this section, we report the results obtained by solving the instances proposed in [35] [^1] Instances available at http://ic.ufal.br/professor/rian/Instances/Instances-MCFST-Zhang.zip., as well as the ones proposed in [6] [^2] Instances available at http://www.dipmat2.unisa.it/people/carrabs/www/DataSet/CMST.zip., with our Kernel Search Method (Algorithms 23). In the following, we refer to the instances proposed in [35] as ZKP instances and to the ones proposed in [6] as CCPR instances. The ZKP instances are generated by using different values for nodes, edges, and number of conflicts and they are classified into two types. By construction, there exists at least one conflict-free solution for all Type 2 instances, while this is not the case for Type 1 instances. In the computational tests, we consider only instances where at least the existence of a feasible solution is guaranteed. The CCPR instances have a number of nodes in $\{25,50,75,100\}$ and a number of edges such that the density of the graph $m(m-1)/2$ is equal to $0.2$, $0.3$, or $0.4$. The number of conflicting pairs $|C|$ is equal to 1%, 4%, 7% of the graph density. For each combination of these parameters (scenario), the authors generated 5 instances by varying the seed used for the generation of random numbers, obtaining a total of 180 instances.
We conducted an extensive parameter tuning process for the algorithm, exploring a range of values for key parameters, and selecting the ones yielding better results. In particular, we set a time limit of 3600 seconds for the whole algorithm and a time limit for each subproblem of 350 seconds for ZKP instances and 420 seconds for CCPR instances. The computing time may exceed the time limit of at most 350 seconds for the ZKP and 420 seconds for the CCPR instances, if the solution of the last subproblem starts after 3250 or 3180 seconds, respectively.
Furthermore, in our tests, we set the starting values of $K,$ and $d$:
$K=\lfloor 1.2(|V|-1)\rceil\quad d=\lfloor 0.1(|E|-K)\rceil,$
which are then updated during the initialization phase, in such a way that the starting kernel set contains a sufficient number of variables to eventually provide a feasible solution to the problem (i.e., $n-1$) plus possible alternatives, and the subproblems solved at each iteration over the kernel set and the buckets are not excessively large. Instead, the maximum number of outer iterations $P$ is set to $1$, for the ZKP instances, and to $4$, for the CCPR instances. Furthermore, we set the parameter $\delta$, indicating the maximum number of iterations the KS can perform without improving, to $\lfloor 0.3\cdot b\rfloor$ (it changes at every iteration $p$, since $b$ changes along these iterations).
@@
Tables 1-2 reports the results obtained by testing the KS method on Type 1 and Type 2 ZKP instances (respectively) proposed in [35] (for which at least a feasible solution is found), compared to the ones obtained by using the following heuristic approaches:
@@
  • the metaheuristic introduced in [14], GRASP-AM which we also use within our Kernel Search initialization phase;
@@
  • the classic Kernel Search algorithm (ClassicKS) as introduced in [1] – never used to solve the MSTC problem – which uses neither the independent set nor the GRASP. It performs at most $P=4$ iterations over the buckets for both types of instances, merging them two by two at each iteration. The initial kernel set size is $K$, and the initial bucket size is $d$. The stopping criterion is the same used in our KS algorithm, with $\delta=\lfloor 0.3\cdot b\rfloor$.
Each subproblem involved in the KS and ClassicKS algorithms is solved using the state-of-the-art solver Cplex. They are coded in Python on an OSX platform, running on an Intel(R) Core(TM) i7-10700K CPU 3.80GHz with 32 GB of RAM, equipped with the IBM ILOG CPLEX (Version identifier: 22.1.0.0) solver (multi-thread mode). As concerns the time, in order to have a fair comparison, the CPU time reported in the literature for each method is properly scaled according to the Whetstone benchmark [33].
@@
ID $n$ $m$ $|C|$ UB_best time gap time gap time gap time gap time gap time gap
1 50 200 199 $*{}^{*}$ 0.46 0.00 0.22 2.54 0.13 0.00 0.63 0.00 0.57 0.00 0.54 0.00
2 50 200 398 $*{}^{*}$ 0.44 0.00 0.29 1.30 0.13 0.00 0.79 0.00 16.42 0.00 0.85 0.00
3 50 200 597 $*{}^{*}$ 0.40 0.00 0.36 28.24 0.13 0.00 1.12 0.00 9.33 3.82 0.84 0.00
4 50 200 995 $*{}^{*}$ 0.42 0.91 0.51 19.86 0.13 0.00 1.68 0.00 1.49 30.06 0.67 0.00
5 100 300 448 $*{}^{*}$ 1.54 1.16 0.76 1.44 1.34 0.00 4.80 0.00 135.72 0.00 15.04 0.00
6 100 300 897 $*{}^{*}$ 1.16 7.72 1.38 100 1.21 0.00 5.27 0.07 377.76 0.67 4.72 0.00
7 100 500 1247 $*{}^{*}$ 3.33 0.00 1.90 0.61 1.30 0.00 4.68 0.00 207.63 0.16 9.13 0.00
8 100 500 2495 $*{}^{*}$ 3.29 3.37 3.53 36.97 3.22 0.00 10.81 0.00 1016.28 1.62 10.90 0.00
9 100 500 3741 7523 2.39 1.89 4.38 100 3.06 0.00 15.79 0.00 3944.75 100 5.18 0.00
13 200 600 1797 13831 46.58 8.60 5.51 100 23.18 1.67 42.35 0.00 3881.71 100 38.02 -0.06
16 300 800 3196 21480 67.62 2.75 12.17 100 38.68 0.89 29.36 0.00 1591.68 8.20 54.47 -0.18
Average values 11.60 2.40 2.82 44.63 6.59 0.23 10.66 0.01 1020.30 22.23 12.76 -0.02
%best 36.4 0 81.8 72.7 27.3 100
@@
[…]
The results of the average values line show that all the algorithms but ClassicKS are very fast because, on average, most of them require less than 13 seconds. The fastest algorithm is HDA+, which requires only 2.82 seconds but is the less effective one with an average gap of 44.63%. On the contrary, the most effective algorithm is KS which always finds the optimal solution, in the instances where it is known, and, for the last two instances, it reports a solution better than the best one known in the literature. Its final average gap is equal to -0.02% and these results are obtained, on average, in 12.76 seconds with a peak of 54 seconds. The effectiveness of KS in these instances is further certified by %best line where its value is equal to 100% since it always finds a solution that is at least as good as the best one known in the literature. Notice that it is the only algorithm to obtain this result. Similar results are obtained by TPMT and GRASP-AM which are faster than KS but show a higher average gap. Moreover, their %best value is equal to 72.7% and 81.8%, respectively. It is worth noting the huge difference between the results of ClassicKS and KS algorithms. ClassicKS is the slowest algorithm of the set and only in three cases, it shows a gap equal to 0%, while its average gap is very high (22.23%). These preliminary results already show the impact that the use of the independent set strategy and the GRASP algorithm have on the effectiveness and performance of ClassicKS.
@@
As far as it concerns Type 2 instances, in the literature they are solved by MEGA, and TPMT algorithms only. For this reason, Table 2 reports the results of these two algorithms as well as ClassicKS and KS. This set of instances appears very easy to solve for all the algorithms because all of them find the optimal known solution in less than 10 seconds, on average. Regarding the performance, the fastest algorithm is TPMT, which requires always less than 4 seconds, followed by MEGA with a peak of 25.59 seconds. Finally, ClassicKS and KS are slower with a peak of around 1 minute.
Figure 1: Performance comparison among HDA+, ClassicKS and KS algorithms on the CCPR instances.
@@
The results on the CCPR instances are reported in Table 3 in the Appendix. In this table, we only compare our approach with HDA+ method, besides ClassicKS, because the other algorithms were not tested on them. As Table 3 contains the results of 151 instances, we decided to move it into the Appendix and to represent here its results by using the cumulative charts reported in Figures 1 and 2.
@@
The horizontal axis in Figure 1 reports the computational time in seconds, and the vertical one shows the percentage of best/optimal solutions found within that time. More precisely, a $(x,y)$ point on this plot means that within $x$ seconds of computation, the algorithm found the best/optimal solution for yof the instances in at most $x$ seconds. This implies that the faster the growth of a curve, the better the performance. It is worth noting that in our tables we compare the solutions of our algorithms with the best ones available in the literature (i.e. UB_best value). As a consequence whenever KS finds a solution which is better than UB_best, the corresponding percentage gap will be negative. Since negative percentages are not allowed in the graphic of Figure 1 then the comparison is not carried out with UB_Best, but with the best solution obtained by all the algorithms of the literature as well as our algorithms. We refer to this new value as New_UB_Best, and we will use it also for the construction of the graphic in Figure 2. The blue curve in Figure 1 is associated with the KS algorithm, the red one corresponds to ClassicKS, and the green one to the HDA+ algorithm. Figure 1 certifies the superiority of the KS with respect to the other two algorithms. Indeed, KS reaches the $\sim$68of best/optimal solutions found, while for ClassicKS and HDA+ this percentage is equal to $\sim$46and $\sim$30only, respectively. By comparing the ClassicKS and the KS curves, we observe that ClassicKS requires 262 seconds to reach its peak of $\sim$46whereas KS reaches this percentage in only 25 seconds. Moreover, after 262 seconds, KS finds the $\sim$58of best/optimal solutions which represents a relevant improvement with respect to ClassicKS. HDA+ requires only 2 seconds to reach its peak of $\sim$30but KS obtains a very close percentage (around 29%) at the same time.
Figure 2 shows the effectiveness of the three algorithms in the 151 instances considered. The horizontal axis reports the percentage gap from New_UB_Best at the end of the computation, whereas the vertical one shows the percentage of instances for which the percentage gap returned by the algorithm at termination is lower than or equal to the value reported on the horizontal axis. As for the graphic in Figure 1, the faster the growth of a curve, the better the effectiveness of the corresponding algorithm.
It is easy to see from this graphic that KS overcomes the other two algorithms because its curve is always higher than the other two ones. This means that KS obtains a lower percentage gap from the best solution, for more instances with respect to the other two algorithms.
@@
When the horizontal value is 0%, the curves of Figure 2 show the percentage of instances for which the solution of the algorithm matches the New_UB_Best value, and this percentage is equal to $\sim$68%, $\sim$46and $\sim$30for KS, ClassicKS and HDA+, respectively. These percentages clearly coincide with the peaks reported in Figure 1 for these three algorithms. The percentage of instances for which the gap is lower than 5% significantly increases for all the algorithms. This percentage is, indeed, equal to $\sim$96%, $\sim$79and $\sim$61for KS, ClassicKS and HDA+, respectively. These results highlight the effectiveness of KS which almost always returns a solution very close to the best one and obtains a gap lower than or equal to 10% for the 98% of instances. Within the same threshold of 10%, the percentage of ClassicKS and HDA+ equals $\sim$79and $\sim$67%, respectively, which are significantly lower than the one of KS.

### v2

@@ 6 Computational Tests
In this section, we report the results obtained by solving the instances proposed in [38] [^1] Instances available at http://ic.ufal.br/professor/rian/Instances/Instances-MCFST-Zhang.zip., as well as the ones proposed in [7] [^2] Instances available at http://www.dipmat2.unisa.it/people/carrabs/www/DataSet/CMST.zip., with KS (Algorithms 34). In the following, we refer to the instances proposed in [38] as ZKP instances and to the ones proposed in [7] as CCPR instances. The ZKP instances are generated by using different values for nodes, edges, and number of conflicts and they are classified into two types. By construction, there exists at least one conflict-free solution for all Type 2 instances, while this is not the case for Type 1 instances. In the computational tests, as for the Type 1, we consider only the 11 instances where at least the existence of a feasible solution is guaranteed. The CCPR instances have a number of nodes in $\{25,50,75,100\}$ and a number of edges such that the density of the graph $m(m-1)/2$ is equal to $0.2$, $0.3$, or $0.4$. The number of conflicting pairs $|C|$ is equal to 1%, 4%, 7% of the graph density. For each combination of these parameters (scenario), the authors generated 5 instances by varying the seed used for the generation of random numbers, obtaining a total of 180 instances.
6.1 Parameter tuning
In this subsection, we explore the impact of various parameter settings on KS performance across different instances. We conducted an extensive parameter tuning process, exploring a range of values for the key parameters: $K$, $d$, $\Delta$, $P$. The goal is to identify the configurations that yield the best results in terms of both output quality and execution time. To conduct a comprehensive analysis, we set up a series of experiments where we systematically varied key parameters of the algorithm. Each configuration is evaluated on multiple instances, allowing us to observe how different parameter settings influence the performance.
For the heuristic procedure, based on preliminary experiments, we set $\bar{h}=20$ and $\bar{t}=500$. For the whole algorithm, we set a time limit of $3600$ seconds and we test a time limit for each subproblem of $180,300,420$ seconds.
As for the dimension of the starting kernel set $K$ and the dimension of the starting buckets $d$, we need to take into account the dimension of the graph, in such a way that the starting kernel set contains a sufficient number of variables to eventually provide a feasible solution to the problem (i.e., $n-1$) plus possible alternatives, and the subproblems solved at each iteration over the kernel set and the buckets are not excessively large. For this reason, we considered the following formulas
$K=\lfloor\alpha\cdot(|V|-1)\rceil\quad d=\lfloor\beta\cdot(|E|-K)\rceil,$
with $\alpha$ varying in $\{1.05,1.1,1.15,1.2,1.25\}$ and $\beta$ in $\{0.15,0.2,0.25\}$. Instead, the maximum number of outer iterations $P$ is set to $2,3,4$. As for the parameter $\Delta$, indicating the maximum number of iterations KS can perform without improving, we need to link it to the number of buckets, that is, $b$. We set it to $\lfloor\delta\cdot b\rfloor$, with $\delta\in\{0.3,0.4,0.5,0.6\}$. We remark that $\Delta$ changes at every iteration $p$, since $b$ changes along these iterations.
In our preliminary experiments, we considered 9 Type 1 ZKP instances with up to 100 nodes and 500 edges and 10 CCPR instances with 50 nodes, 245 edges, and up to 1196 conflicts. We report, in Tables 2, for each parameter value the percentage of instances solved to optimality together with the average running time. For the final tests, we select the configuration associated with the maximum weighted average between the percentage of instances for which we find a solution that is at least as good as the best found in the literature (with weight $0.8$) and the inverse of the average running time (with weight $0.2$). In the case of ties, we select the configuration with a greater percentage of optimally solved instances and then the one with a shorter time. For the ZKP instances, the selected parameter values are $P=4,\alpha=1.1,\beta=0.2,\delta=0.6,$ and the inner time limit is $420$ seconds. For the CCPR instances, the selected parameter values are $P=4,\alpha=1.2,\beta=0.2,\delta=0.4,$ and the inner time limit is $180$ seconds.
(a) ZKP instances
parameter % solved instances average time[s] weighted average
$P$
2 36.05 7.90 0.16
3 47.84 14.90 0.20
4 53.15 29.89 0.22
$\alpha$
1.05 40.12 19.69 0.17
1.10 49.79 29.94 0.20
1.15 49.59 20.55 0.20
1.20 41.67 7.26 0.18
1.25 47.22 14.99 0.20
$\beta$
0.15 46.11 19.89 0.19
0.20 52.10 17.10 0.21
0.25 38.83 20.05 0.16
$\delta$
0.30 45.43 20.51 0.19
0.40 45.51 16.35 0.19
0.50 45.84 17.90 0.19
0.60 45.93 20.72 0.19
timelimit
180 45.31 14.83 0.19
300 45.62 18.21 0.19
420 46.11 23.50 0.19
(b) CCPR instances
parameter % solved instances average time[s] weighted average
$P$
2 61.17 2.26 0.29
3 66.00 2.69 0.30
4 66.00 2.68 0.30
$\alpha$
1.05 66.67 4.51 0.29
1.10 63.33 3.57 0.28
1.15 66.39 2.18 0.31
1.20 65.56 0.81 0.39
1.25 66.00 1.61 0.30
$\beta$
0.15 65.17 4.82 0.28
0.20 64.00 1.31 0.33
0.25 64.00 1.48 0.32
$\delta$
0.30 64.22 2.48 0.30
0.40 64.44 2.56 0.30
0.50 64.44 2.57 0.30
0.60 64.44 2.58 0.30
timelimit
180 64.39 2.50 0.30
300 64.39 2.52 0.30
420 64.39 2.51 0.30
Table 2: Tuning of the parameters.
6.2 Comparisons with existing algorithms
@@
In this section, we report the results obtained by testing our KS, in comparison with state-of-the-art approaches. Tables 3-4 report the results on Type 1 and Type 2 ZKP instances (respectively) proposed in [38] (for which at least a feasible solution is found), compared to the ones obtained by using the following heuristic approaches:
@@
  • the metaheuristic introduced in [15], GRASP-AM, which uses a GRASP enhanced with an adaptive memory strategy to efficiently explore the solution space and improve solution quality;
@@
  • the classic KS algorithm (ClassicKS) as introduced in [1] – never used to solve the MSTC problem – which uses neither the independent set nor the heuristic. It performs at most $P$ iterations over the buckets for both types of instances, merging them two by two at each iteration. The stopping criterion is the same used in our KS algorithm, and we set $P=4,\alpha=1.2$, $\beta=0.1$, $\delta=0.3$, and the inner time limit to 420 seconds. These values are based on some preliminary experiments performed on a small subset of instances.
Following the approach used by MEGA [6], GRASP-AM [15], TMPT [12] our proposed KS is executed for different runs on each instance (5 for the ZKP and 10 for the CCPR instances), each for a certain seed, which only affects the result found by the StartingSolution heuristic, used within the Initialization Phase. Each subproblem involved in KS and ClassicKS algorithms is solved using the state-of-the-art solver Cplex. They are coded in Python on an OSX platform, running on an Intel(R) Core(TM) i7-10700K CPU 3.80GHz with 32 GB of RAM, equipped with the IBM ILOG CPLEX (Version identifier: 22.1.0) solver. As concerns the time, in order to have a fair comparison, the CPU time reported in the literature for each method is properly scaled according to the Whetstone benchmark [36].
The heading of Tables 3-4 are the following: ID is the identifier of each instance; $n$ is the number of nodes of the graph; $m$ is the number of edges of the graph; $|C|$, is the number of conflicts; UB_best is the best value found either by the state-of-the-art methods we use to benchmark our approach or by the exact Branch and Cut method proposed in [7]; time, the minimum computational time (in seconds) required by the method to return its best solution value; gap which is the percentage gap of the best upper-bound $UB$ found over the multiple runs with respect to UB_best, i.e., gap $=\frac{100(\text{{UB}}-\text{{UB\_best}})}{\text{{UB\_best}}}.$ Whenever UB_best is certified to be optimal by the Branch and Cut, we report its value with the “*” symbol. Only in Table 3, for KS, we also report, in the UB_time column, the time needed by the algorithm to obtain the corresponding $UB$. We do not report this information in Table 4 as UB_time corresponds to the whole time for the ZKP Type 2 instances. All the instances we solve are preprocessed according to the procedure described in 5.3. The time needed for this preprocessing as well as the time needed for all the algorithms involved within KS are included in the reported computational time. At the bottom of each table, the row “Average values” shows the average values for each column, and the row “%best” reports the percentage of instances having a gap lower than or equal to 0%. We highlight in bold the best average time and the best average gap.
@@
ID $n$ $m$ $|C|$ UB_best time gap time gap time gap time gap time gap time UB_time gap
1 50 200 199 708 0.46 0.00 0.22 2.54 0.13 0.00 0.63 0.00 0.57 0.00 2.90 1.38 0.00
2 50 200 398 770 0.44 0.00 0.29 1.30 0.13 0.00 0.79 0.00 16.42 0.00 1.24 0.57 0.00
3 50 200 597 917 0.40 0.00 0.36 28.24 0.13 0.00 1.12 0.00 9.33 3.82 2.13 1.84 0.00
4 50 200 995 1324 0.42 0.91 0.51 19.86 0.13 0.00 1.68 0.00 1.49 30.06 3.33 1.07 0.00
5 100 300 448 4041 1.54 1.16 0.76 1.44 1.34 0.00 4.80 0.00 135.72 0.00 5.16 3.90 0.00
6 100 300 897 5658 1.16 7.72 1.38 100 1.21 0.00 5.27 0.07 377.76 0.67 55.18 18.80 0.00
7 100 500 1247 4275 3.33 0.00 1.90 0.61 1.30 0.00 4.68 0.00 207.63 0.16 10.70 3.32 0.00
8 100 500 2495 5997 3.29 3.37 3.53 36.97 3.22 0.00 10.81 0.00 1016.28 1.62 75.45 26.35 0.00
9 100 500 3741 7523 2.39 1.89 4.38 100 3.06 0.00 15.79 0.00 3985 100 308.82 24.77 0.00
13 200 600 1797 13839 46.58 8.60 5.51 100 23.18 1.67 42.35 0.00 3881.71 100 1726.48 522.18 -0.06
16 200 800 3196 21518 67.62 2.75 12.17 100 38.68 0.89 29.36 0.00 1591.68 8.20 631.45 229.59 -0.18
Average values 11.60 2.40 2.82 44.63 6.59 0.23 10.66 0.01 1020.30 22.23 256.62 75.80 -0.02
%best 36.4 0 81.8 90.9 27.3 100
@@
As far as it concerns Type 1 instances (Table 3) the results of the average values line show that the fastest algorithm is HDA+, which requires only 2.82 seconds but is the less effective one with an average gap of 44.63%. On the contrary, the most effective algorithm is KS which always finds the optimal solution, in the instances where it is known, and, for the last two instances, it reports a solution better than the best one known in the literature. Its final average gap is equal to -0.02% and these results are obtained, on average, in around 75 seconds.
The effectiveness of KS on these instances is further certified by %best line where its value is equal to 100% since it always finds a solution that is at least as good as the best one known in the literature. Notice that it is the only algorithm to obtain this result. Similar results are obtained by GRASP-AM and TPMT which are faster than KS but show a higher average gap. Moreover, their %best value is equal to 81.8% and 90.91%, respectively. It is worth noting the big difference between the results of ClassicKS and KS. ClassicKS is the slowest algorithm of the set and only in three cases it shows a gap equal to 0%, while its average gap is very high (22.23%). These preliminary results already show the impact that the use of the independent set strategy and the heuristic algorithm has on the effectiveness and performance of the KS method.
@@
As regards Type 2 instances (Table 4), in the literature they are solved by MEGA, and TPMT algorithms only. For this reason, Table 4 reports the results of these two algorithms as well as ClassicKS and KS. This set of instances appears very easy to solve for all the algorithms because all of them find the optimal known solution in less than 0.10 seconds, on average.
The results on the CCPR instances are reported in Table LABEL:tab:kernel_cmst in the Appendix. In this table, we only compare KS with HDA+ method, besides ClassicKS, because the other algorithms were not tested on them. As Table LABEL:tab:kernel_cmst contains the results of 151 instances, we decided to move it into the Appendix and to represent its results both in Table 5 by aggregating them for the dimension of the instances, and in the cumulative charts reported in Figures 1 and 2.
Instance HDA+ ClassicKS KS
$n$ $m$ UB_best time gap time gap time UB_time gap
25 60 384.8 0.04 0.25 0.39 0.23 0.31 0.11 0.07
25 90 329.9 0.07 0.29 1.60 0.43 0.77 0.26 0.00
25 120 308.3 0.10 0.84 0.65 0.51 1.33 0.33 0.04
50 245 704.6 0.61 34.9 616.69 8.12 73.79 11.8 -0.39
50 367 644.1 1.47 30.5 727.00 3.87 222.47 26.86 -0.20
50 490 600.8 2.87 26.54 800.76 3.48 309.19 42.17 -0.32
75 555 943.7 3.78 33.28 1189.01 5.28 609.41 87.99 -0.04
75 832 862.9 9.67 10.61 1572.92 0.83 707.01 107.55 0.07
75 1110 862.4 21.92 19.68 1797.31 10.18 867.27 85.20 10.75
100 990 1285.7 17.41 51.46 2001.53 50.02 1084.52 113.25 -3.37
100 1485 1197.7 48.19 51.43 1988.07 50.13 1182.62 93.19 9.16
100 1980 1124.4 102.92 39.53 2004.51 50.00 1207.13 95.62 9.71
Average values 14.16 23.05 924.09 12.72 440.83 47.26 1.73
%best 29.80 47.02 74.83
Table 5: Comparison of the aggregated results on CCPR instances proposed in [7].
Already by looking at the aggregated results in Table 5 it is evident that, despite requiring more computational time w.r.t. HDA+, KS is by far the best in terms of the gap. Specifically, KS achieves the best average gap across all instance groups and the highest percentage of instances with a non-positive gap relative to the best solution value found in the literature (last row of the table). Notably, 74.83% of the CCPR instances are solved with a non-positive gap (15 new best-known solution), the average gap is 1.73% and the average computing time is approximately 7 minutes. Although this is significantly higher than the average computing time of the HDA+ method, the latter yields a much larger average gap of 23.05%. The classic KS approach is slower than our KS, returning an average gap of 12.72%, which is still less than the one returned by the HDA+ method.
Figure 1: Performance comparison among HDA+, ClassicKS and KS on the CCPR instances.
@@
In Figure 1, the horizontal axis reports the computational time in seconds, and the vertical one shows the percentage of best/optimal solutions found within that time. More precisely, a $(x,y)$ point on this plot means that within $x$ seconds of computation, the algorithm found the best/optimal solution for yof the instances in at most $x$ seconds. This implies that the faster the growth of a curve, the better the performance. It is worth noting that in our tables we compare the solutions of our algorithms with the best ones available in the literature (i.e. UB_best value). As a consequence whenever KS finds a solution which is better than UB_best, the corresponding percentage gap will be negative. Since negative percentages are not allowed in the graphic of Figure 1 then the comparison is not carried out with UB_Best but with the best solution obtained by all the algorithms of the literature as well as our algorithms. We refer to this new value as New_UB_Best, and we will use it also for the construction of the graphic in Figure 2. The blue curve in Figure 1 is associated with KS, the pink one corresponds to ClassicKS, and the green one to the HDA+ algorithm. Figure 1 certifies the superiority of KS with respect to the other two algorithms. Indeed, KS reaches the $\sim$75of best/optimal solutions found, while for ClassicKS and HDA+ this percentage is equal to $\sim$47and $\sim$30only, respectively. By comparing the ClassicKS and KS curves, we observe that ClassicKS requires the whole time limit to reach its peak of $\sim$47whereas KS reaches this percentage in less than 200 seconds. HDA+ in total only manages to find 46 best solutions (see Figure 2) and it does it in 2 seconds (then it always obtains a positive gap at termination, that is why the curve stays constant). KS obtains a very close percentage (the 28%) in the same amount time.
@@
Figure 2 shows the effectiveness of the three algorithms in the 151 instances considered. The horizontal axis reports the percentage gap from New_UB_Best at the end of the computation, whereas the vertical one shows the percentage of instances for which the percentage gap returned by the algorithm at termination is lower than or equal to the value reported on the horizontal axis. As for the graphic in Figure 1, the faster the growth of a curve, the better the effectiveness of the corresponding algorithm. It is easy to see from this graphic that KS overcomes the other two algorithms because its curve is always higher than the other two ones. This means that KS obtains a lower percentage gap from the best solution, for more instances with respect to the other two algorithms. When the horizontal value is 0%, the curves of Figure 2 show the percentage of instances for which the solution of the algorithm matches the New_UB_Best value, and this percentage is equal to $\sim$75%, $\sim$47and $\sim$30for KS, ClassicKS and HDA+, respectively. These percentages clearly coincide with the peaks reported in Figure 1 for these three algorithms. The percentage of instances for which the gap is lower than 10% significantly increases for all the algorithms. This percentage is, indeed, equal to $\sim$90%, $\sim$85and $\sim$66for KS, ClassicKS and HDA+, respectively. These results highlight the effectiveness of KS which almost always returns a solution very close to the best one and obtains a gap lower than or equal to 20% for the 95% of instances. Within the same threshold of 20%, the percentage of ClassicKS and HDA+ equals $\sim$88and $\sim$75%, respectively, which are much lower than the one of KS.
In conclusion, while the average termination time of our KS algorithm is always higher than that of other algorithms (except for the classic KS one, where our enhancements lead to improvements also in terms of time), the time required to find the best-returned incumbent is definitely competitive. Moreover, this best solution is 17 times better than those reported in the literature and at least as good as the best-known solutions in 80% of cases.